# Random Forest for Dynamic Risk Prediction of Recurrent Events: A Pseudo-Observation Approach

### v2

@@ 6 Prediction of COPD Exacerbations
We apply RFRE.PO to predict the probability of remaining exacerbation-free over a subsequent $\tau=180$ day follow-up period, using data from Azithromycin for the Prevention of COPD (Albert and others, 2011). Azithromycin for the Prevention of COPD was a multi-center clinical trial, aiming to estimate the treatment effect 250mg of oral azithromycin had on reducing exacerbations in COPD patients. We pooled data from both arms of the clinical trial, and used treatment assignment as one of many covariates in our prediction model.
@@
Our example is applied to study participants ($n=1035$) who have complete baseline data, as well as longitudinal sleep study variables. Recurrent exacerbations were monitored over the following year. There are 962 covariates measured, including demographic variables, social well-being surveys, sleep quality metrics and clinical characteristics such as forced vital capacity (FVC), and medication use. There is strong correlation between variables as some are derived from each other, making RFRE.PO an attractive alternative to model-based predictions (see Supplemental Figure LABEL:fig:corrHeatMap). Censored longitudinal data is constructed with $\mathcal{T}=\{0,30,60,...,180\}$ and used for all analyses shown in this section.
Baseline information on the time since the most recent exacerbation was not available. Hence in this study we are in a scenario similar to the partial history scenario in Section 5. In this example, we create two partial history variables, $H_{p_{1}}(t)$ and $H_{p_{2}}(t)$. $H_{p_{1}}(t)$ was assumed to take a categorical form based on whether the participant had an exacerbation in the past zero to 30 days, 31-92 days, 93-182 days, 183 days to 365 days, or finally, if that time was more than 365 days or never. $H_{p_{2}}(t)$ is 100 times the estimated 30-day event rate using all follow-up prior to time $t$ (so as to provide point estimates on a reasonable scale). For the first $H_{p_{1}}(0)$ and $H_{p_{2}}(0)$, multiple imputation was based on event rates in the placebo group for all participants, and Rubin’s rule (Rubin, 1987) was used to combine inference across multiply imputed datasets as appropriate.
@@
For comparison purposes, we present RFRE.PO results alongside results from two modified XMT models, where modified XMT models are fit via the geese command from the geepack library in R. The first modified XMT model (Model 1) uses the same predictors identified in Xia, Murray and Tayob (2020) for this cohort, where these predictors were based on clinical judgement for estimating treatment effect, adjusted for known confounders. The second modified XMT model (Model 2) is based on Wald forward selection applied to the first multiply imputed data set with $p<0.05$ required for entry; covariates that introduce model instability are discarded from consideration. Prediction algorithms for RFRE.PO and Model 2 were applied to a training cohort (70% of original cohort) and all algorithms were evaluated via the C-statistic in a validation cohort (remaining 30% of original cohort). For any particular multiply imputed validation dataset, standard errors of the C-statistic are calculated via the bootstrap with $b=100$ bootstrap samples. C-statistics for models involving multiply imputed history variables are combined across multiply imputed datasets using Rubin’s rule. Parameter estimates and standard errors for Model 2, which involves the imputed partial history covariates, are combined across multiply imputed validation datasets via Rubin’s Rule. Permutation tests associated with the RFRE.PO algorithm predictors are calculated within the training dataset, using out-of-bag samples as described in Section 3.
@@
For Model 1, where estimation of treatment effect, and not necessarily prediction was the goal, predictors that were identified as important in Xia, Murray and Tayob (2020) were baseline forced expiratory volume in liters (FEV), age in decades (Age10), gender (0 for female, 1 for male), and smoking status at baseline (Smoker):
$\text{logit}\{E({S}^{\tau}_{i}(t))\}=\beta_{0}+\beta_{1}I(\text{Treatment})_{i}+\beta_{2}(\text{Baseline FEV})_{i}+\beta_{3}(\text{Age10})_{i}+\beta_{4}I(\text{Male})_{i}+\beta_{5}I(\text{Smoker})_{i},$
@@
where parameter estimates for the modified XMT model are displayed in Supplemental Table LABEL:table:summer-estimates.
Model 2, the Wald forward selection model, is:
$\text{logit}\{E({S}^{\tau}_{i}(t))\}=\beta_{0}+\beta_{1}I(H_{p_{1}}=\text{Exacerbation in 0 -30 days prior to time }t)_{i}+$
$\quad\beta_{2}I(\text{Bronchiectasis diagnosis before time }t)_{i}+\beta_{3}I(\text{Male gender})_{i}+\beta_{4}I(\text{Azithromycin group})_{i}+$
$\quad\beta_{5}I(\text{Non-selective beta-blocker use at time }t)_{i}+\beta_{6}I(\text{Mixed beta-blocker use at time }t)_{i}+$
@@
$\quad\beta_{7}(\text{FEV}_{1}\text{/FVC \% predicted at time }t)_{i}+\beta_{8}(\text{SGRQ symptoms at time }t)_{i}+$
$\quad\beta_{9}I(\text{Supplemental oxygen use at time }t)_{i}+\beta_{10}I(\text{Hospitalization one year before study start})_{i}+$
$\quad\beta_{11}H_{p_{2}}(t)_{i}+\beta_{12}I(\text{Inhaled corticosteroid and long-acting muscarinic antagonist use at time }t)_{i}+$
@@
$\quad\beta_{13}I(\text{Inhaled corticosteroid and long-acting beta-agonist use at time }t)_{i},$
@@
For the RFRE.PO algorithm, Supplemental Table LABEL:table:perm-tests displays the corresponding p-values of the permutation tests, combined across multiply imputed datasets via Rubin’s rule. Predictors in Supplemental Table LABEL:table:perm-tests had a statistically significant permutation test z-score for at least one of the multiply imputed training datasets. Additional predictors that did not achieve statistical significance via the permutation test, but were used in the RFRE.PO algorithm in some form, are not shown.
Figure 4 displays the Z-scores for predictors included in Models 1 and 2 (validation dataset) or Table LABEL:table:perm-tests (out-of-bag samples from training dataset). As seen in Section 5, history variables were among the most important variables in the RFRE.PO approaches as well as the modified XMT model selected through Wald Z-score forward selection. While there is overlap between predictors included in the various algorithms, in terms of the C-statistic evaluated in the validation cohort, the RFRE.PO algorithm outperformed both the algorithm based on clinical input (Model 1) and the Wald forward selection model (Model 2) (C-statistic $0.616$ compared to $0.565$ and $0.550$, respectively, see Table 2). A large number of variables that were included in the random forest algorithm had statistical significance at the level of $0.05$ via the permutation test in at least one of the multiply imputed datasets (see Figure 4 for the z-statistic and p-values in Supplemental Table LABEL:table:perm-tests for the predictors deemed important). Statistically significant variables from Model 1 are similarly listed as important RFRE.PO predictors. Variables identified via Wald forward selection in Model 2 are for the most part identified as important variables in the RFRE.PO algorithm. The version of the predictors may vary between these two approaches. For example, variables related to the long-acting muscarinic or bronchodilator, and/or steroid treatment are incorporated in different forms into these models. Similarly, different heart medications are featured in Model 2 (beta-blockers) versus the RFRE.PO algorithm (anti-coagulants). The RFRE.PO algorithm highlights a few variables that were not used in Model 2, including pain level at time $t$ and leukocytosis present at time $t$. The RFRE.PO algorithm has no problem including information from correlated predictors; for instance two included variables measure pain (pain level at time $t$, standardized pain level at time $t$) and two included variables measure general health (general health score at time $t$ standardized general health score at time $t$).
The C-statistic corresponding to Model 1, which benefits from prior clinical knowledge of the causal pathway, validated well. The RFRE.PO random forest algorithm includes predictors, with corresponding split decisions, that minimize squared error of the predictions relative to the observed data. This results in predictor information that is included without rigorous checks for statistical significance, so validation in an independent data set is quite important when using this type of algorithm. Because Model 2 requires statistical significance of all included predictors, its individual predictors already have some theoretical validity built into the selection mechanism. However, with variable power to detect statistical significance of useful predictors, the comparisons with methods that do not require statistical significance will leave Model 2-style strategies at a disadvantage.
@@
The RFRE.PO algorithm is completely agnostic to statistical significance and scientific experience, so long as the squared error within nodes along cut-points decreases. And although it may occasionally include a predictor that is not useful in the validation set (Type I Error analog), it has a much better chance of including predictors that are useful but would not have met statistical significance (akin to a Type II error) in traditional statistical model building environments. If the scientific goal of analysis is prediction, the RFRE.PO approach seems remarkably simple, effective, and reproducible.
Table 2: Comparison of the C-statistics of models fit to the validation set of the Azithromycin in the Prevention of COPD cohort. Models 1 and 2 are fit using the modified XMT approach. Model 1 is the model selected using clinician input from Xia, Murray and Tayob (2020). Model 2 includes predictors through automated forward selection requiring $p<0.05$ for entry into the model. Missing values in Model 2 and RFRE.PO for time since most recent exacerbation are multiply imputed using the coding scheme described in Section 6. RFRE.PO processes our random forest approach using the same predictors considered in Model 2. C-statistic estimates and standard errors for all methods and models are averaged across results from multiply imputed data with Rubin’s rule applied to bootstrapped ($b=100$) C-statistic standard errors.
Model Name C-Statistic
Mean Standard Error
RFRE.PO $0.616$ $0.009$
Model 1 $0.565$ $0.009$
Model 2 $0.550$ $0.010$
Figure 4: Wald Statistic Z-scores derived from the validation set for predictors that were included in Models 1 (modified XMT Xia, Murray, Tayob 2020 Model) and 2 (modified XMT Wald Forward Selection) after adjusting for multiple imputation using Rubin’s Rule, where appropriate, and permutation test Z-scores derived from out-of-bag samples as part of the RFRE.PO algorithm, adjusting for multiple imputation using Rubin’s Rule where appropriate. For the RFRE.PO method, 100 permutations were used in calculating the permutation Z-scores. Note that some Wald forward selection variables lost significance as additional terms were added.
@@ 7 Discussion
In this paper we presented a novel method for dynamically estimating probability of event-free survival during different follow-up windows. Using RFRE.PO, we showed the utility of our method in simulation, comparing its performance to traditional longitudinal model fits with an modified XMT model. In all instances, except for the most highly correlated, RFRE.PO outperformed the main effects modified XMT model that a traditional modeler might first attempt to fit. In instances of high correlation, RFRE.PO and the main effects model competed with the ad hoc nonlinear modified XMT model- that is, the perfect model that exactly reflected how the data were simulated.
An attractive feature of random forest applied to longitudinal follow-up windows is that a split based on the follow-up window times can occur and predictors that are more helpful in later follow up times can emerge naturally in importance. We explored recent history prior to the beginning of each follow-up window as a covariate in the algorithm, and found that it was one of the most useful predictors in settings with high correlation between recurrent events. The use of complex scoring algorithms has increased with the development of machine learning predictors such as RFRE.PO. For example, the COPD research community has agreed that algorithmic prediction models need to be user-friendly to be used in clinical practice. To this end, the COPDGene Study ® has created an app to calculate various risk scores based on patient characteristics that a user can input. Apps that simplify risk score calculations along these lines are one such way to bring complicated scores to clinical practitioners.
It is always important when introducing a new piece of methodology to clearly identify research questions that it can and cannot address. Random forest technology in general is aimed towards prediction, and for RFRE.PO, predicting the probability of remaining event-free during subsequent follow-up windows after patient check-ins. In this case, all available predictors that the treating clinician can access are valid and important to use. There is no goal, per sé, of establishing which predictors are statistically significant or have causal relationships with the outcomes. It is likely that some, if not most, of the predictors with high variable importance scores have these features, but metrics assessing statistical significance and causality are not necessarily valid for these algorithms.
In contrast, models such as the modified XMT model for recurrent event outcomes often have the goal of measuring the impact of a particular predictor or set of predictors on the outcome via an effect size, confidence interval and p-value. In the Azithromycin for Prevention of COPD cohort, the azithromycin effect was the key research question, requiring a measure of statistical significance as part of the drug approval process. Some predictors that would be very helpful in prediction are not allowed in these models assessing treatment effect. For instance, it is well-known that one should not include updates to patient characteristics after treatment has been initiated if these characteristics could have been improved by treatment; doing so would adjust away part of the treatment effect under study. Similarly, clinical trial analyses often require careful adjustment for confounders. If the goal is prediction and not assessment of treatment effect, these same confounders might not be included in the algorithm and in fact may not be helpful relative to other available predictors.
The ability of flexible algorithms like RFRE.PO to sift through large numbers of correlated predictors with complicated relationships to outcomes is an argument for their use over traditional reliance on models such as the modified XMT model. The growing popularity of machine learning tools is likely to dominate analyses of high-dimensional data, and the RFRE.PO algorithm contributes to this literature as one of the first machine learning methods to be able to make dynamic predictions from recurrent event data.
@@ Acknowledgment
The research is supported in part by a Precision Health Investigator Award from University of Michigan, Ann Arbor (AL, ZW). Conflict of Interest: None.

### v3

@@ 8 Prediction of COPD Exacerbations
Azithromycin for the Prevention of COPD (Albert and others, 2011) was a multi-center clinical trial evaluating the effect of daily oral azithromycin (250mg) in preventing severe COPD exacerbations during a year of follow-up. We apply RFRE.PO to predict the probability of remaining exacerbation-free over a subsequent $\tau=180$ day follow-up window in this cohort ($n=1035$ study participants with complete baseline data and longitudinal sleep study variables). There are 962 available covariates, including demographic variables, social well-being surveys, sleep quality metrics and clinical characteristics such as forced vital capacity (FVC), and medication use. There is strong correlation between variables as some are derived from each other, making RFRE.PO an attractive alternative to model-based predictions (see Supplemental Figure LABEL:fig:corrHeatMap). Censored longitudinal data is constructed with $\mathcal{T}=\{0,30,60,...,180\}$, and used for all analyses, with $T_{i}=\{0,30,\ldots t_{b_{i}}\}$ for individual $i$.
Baseline information on recurrent event times prior to $t=0$ were not available. We defined a full history covariate, $H_{f,i}(t)$, as $I$(Severe exacerbation recorded prior to $t$), with $H_{f,i}(0)$=0. We also considered two partial covariate history variables from Section 2, the categorical version of $H_{p,i}^{(2)}(t)$ given in Equation 3, with $H_{p,i}^{\dagger}(t)$ replacing $H_{f,i}^{\dagger}(t)$, and $H_{p,i}^{(3)}(t)$. For categorical $H_{p,i}^{(2)}(t)$, we assume $\tau^{*}=1$ year and $\mathcal{T}^{*}_{i}(t)=\{t^{*}\in\mathcal{T}_{i}\cup-1\mbox{ year}:t^{*}+1\mbox{ year}\leq t\}$. This variable categorizes $H_{p,i}^{\dagger}(t)$ in the same manner Equation 3 categorizes $H_{f,i}^{\dagger}(t)$ with higher values indicating shorter exacerbation-free periods leading up to $t$. For $H_{p,i}^{(3)}(t)$, we assume $\tau^{*}=30$ days and $\mathcal{T}^{*}_{i}(t)=\{\mathcal{T}_{i}\cup-30\mbox{ days}:t^{*}+30\mbox{ days }\leq t\}$, so that this variable estimates the probability that prior follow-up windows had an exacerbation within 30 days. Imputed event times prior to time 00 were sampled from an exponential($\lambda_{i}$) distribution with $1/\lambda_{i}$ generated from a uniform(100 day, 180 day) distribution, which was comparable to placebo group outcomes. Rubin’s rule (Rubin, 1987) was used to combine inference across multiply imputed datasets as described in Section 2.
@@
For comparison purposes, we present RFRE.PO results alongside results from two m-XMT models. The first m-XMT model (Clinical Input) includes baseline variables used by Xia, Murray and Tayob (2020) for this cohort; these predictors are commonly used to assess confounder-adjusted treatment effects in COPD journals. The second m-XMT (Forward Selection) model is based on Wald forward selection applied to the first multiply imputed data set with $p<0.05$ required for entry; covariates that introduce model instability are discarded from consideration. Prediction algorithms for RFRE.PO and the forward selection model were applied to a training cohort (70% of original cohort) and all algorithms were evaluated via the C-statistic in a validation cohort (remaining 30% of original cohort). For any particular multiply imputed validation dataset, standard errors of the C-statistic are calculated via the bootstrap with $b=100$ bootstrap samples. C-statistics for models involving multiply imputed history variables are combined across multiply imputed datasets using Rubin’s rule. Parameter estimates and standard errors for m-XMT (Forward Selection), which involves the imputed partial history covariates, are combined across multiply imputed validation datasets via Rubin’s Rule. Permutation tests associated with the RFRE.PO algorithm predictors are calculated within the training dataset, using out-of-bag samples as described in Section 5.
@@
For m-XMT (Clinical Input), where estimation of treatment effect, and not necessarily prediction was the goal, predictors are baseline forced expiratory volume in one second in liters (Baseline FEV1), age in decades (Age10), male gender {$I$(Male)}, and baseline smoking status {$I$(Smoker)}:
[…]
@@
where parameter estimates for the m-XMT model are displayed in Supplemental Table LABEL:tab:summer-estimates.
The m-XMT model based on Wald forward selection, is:
$\text{logit}\{E({S}^{\tau}_{i}(t))\}=\beta_{0}+\beta_{1}I\{\text{Categorical }H_{p,i}^{(2)}(t)=0\}+\beta_{2}I(\text{Hospitalization in year before $t=0$})_{i}$
$\quad+\beta_{3}I(\text{Corticosteroid use in year before $t=0$})_{i}+\beta_{4}H_{f,i}(t)+\beta_{5}t+\beta_{6}I(\text{Male})_{i}+\beta_{7}I(\text{Race = Black})_{i}$
[…]
$\quad+\beta_{10}I(\text{Seen a mental health provider at time }t)_{i}+\beta_{11}I(\text{Has pain at time }t)_{i}$
$\quad+\beta_{12}(\text{St. George Respiratory Questionnaire symptom score at time }t)_{i}+\beta_{13}I(\text{Antiarrhythmic use at time }t)_{i}$
@@
$\quad+\beta_{14}(\text{FEV}_{1}\text{/FVC \% predicted at time }t)_{i}+\beta_{15}I(\text{Azithromycin treatment group})_{i}$
@@
$\quad+\beta_{16}I(\text{Inhaled corticosteriod or long-acting beta-agonist use at time }t)_{i}+\beta_{17}I(\text{Other baseline medications})_{i}$
$\quad+\beta_{18}(\text{Supplemental oxygen use at time }t)_{i}+\beta_{19}I(\text{Mixed beta-blocker use at time }t)_{i}$
$\quad+\beta_{20}I(\text{Bronchiectasis diagnosis at time }t)_{i}$
@@
For the RFRE.PO algorithm, Supplemental Table LABEL:table:perm-tests highlights predictors that had a statistically significant permutation test Z-score for at least one of the multiply imputed training datasets. Additional predictors that did not achieve statistical significance via the permutation test, but were used in the RFRE.PO algorithm in some form, are not shown.
Figure 4 displays validation dataset Z-scores for predictors included in the m-XMT (Clinical Input) and m-XMT (Forward Selection) algorithms as well as out-of-bag sample Z-scores for RFRE.PO predictors highlighted in Supplemental Table LABEL:table:perm-tests. In algorithms that included history covariates {m-XMT (Forward Selection) and RFRE.PO}, these predictors showed strong validation results in Figure 4. When prediction algorithms were applied to individuals in the validation cohort, the RFRE.PO algorithm outperformed both the m-XMT (Clinical Input) and m-XMT (Forward Selection) algorithms in terms of both the C-statistic (C-statistic $0.613$ compared to $0.564$ and $0.571$, respectively) and estimated mean squared-error (MSE $0.242$ compared to $0.252$ and 0.264)0.264)0.264 ); See Table LABEL:table:data-res. Time-dependent C-statistics and $MSE(t)$ for follow-up windows starting at time, $t$, in the validation dataset are displayed in Figure LABEL:fig:TimeVaryingCMSE for the various prediction algorithms. Compared to the other algorithms, RFRE.PO has fairly stable and attractive performance metrics across the various follow-up windows, likely because interactions with the various predictors over time are naturally incorporated if useful for prediction.
As can be seen in Figure 4, variables identified via m-XMT models are for the most part identified as important variables in the RFRE.PO algorithm, although the version of the predictors may vary. For example, variables related to the long-acting muscarinic or beta-blockers, and/or steroid treatment are incorporated in different forms into these models (highlighted on the y-axis in purple). Similarly, different variables relating to patient well-being (highlighted in orange) are featured in the m-XMT (Forward Selection) model (has pain at time $t$) versus the RFRE.PO algorithm (pain score at time $t$ and standardized pain score at time $t$), as well as different measures of pulmonary function (highlighted in pink). The RFRE.PO algorithm highlights a few variables that were not used in the forward selection model, including general health score at time $t$ and leukocytosis present at time $t$. The RFRE.PO algorithm has no problem including information from correlated predictors; for instance RFRE.PO included all history variables, whereas this introduced model instability in the m-XMT (Forward Selection) model. Figure 5 displays a $3\times 3$ panel heat map of survival probabilities with corresponding history covariates. Within a row of panels, the general scheme of the predicted values are reflected in the values of derived history covariates, indicating the utility of these covariates in creating dynamic predictions for the Azithromycin for the Prevention of COPD cohort.
@@
The RFRE.PO algorithm includes predictors, with corresponding split decisions, that minimize squared error of the predictions relative to the pseudo-observations. This algorithm is completely agnostic to statistical significance and scientific experience, so long as the squared error within nodes along cut-points decreases, so validation checks are quite important. With variable power to detect statistical significance of useful predictors, methods that require statistical significance for model inclusion may be disadvantaged in terms of prediction when compared to the RFRE.PO algorithm that does not have this requirement. And although the RFRE.PO algorithm may occasionally include a predictor that is not useful in the validation set (type I error analog), it has a much better chance of including predictors that are useful but would not have met statistical significance (akin to a type II error) in traditional statistical model building environments. If the scientific goal of analysis is prediction, the RFRE.PO approach seems remarkably simple and effective.
@@ 9 Discussion
We presented two novel approaches, RFRE.PO and the m-XMT model, for dynamically estimating probabilities of being event-free across a sequence of follow-up windows. An additional novelty is our advice for incorporating covariate history into prediction algorithms, with a multiple imputation approach when covariate history is not known prior to time $t=0$. For purposes of prediction, RFRE.PO tends to outperform the m-XMT (Main Effects) model that a traditional modeler might first attempt. However, the m-XMT model has the advantage of interpretable parameters and a clear formula for prediction that offers more transparency to users, and may accommodate further interactions than explored in the m-XMT (Main Effects) model.
Our presentation did not touch on how causal relationships might be evaluated using these methods. In the Azithromycin for Prevention of COPD study, assessment of the treatment effect disallows using information on time-dependent predictors that could have been changed by treatment after time $t=0$. To do so would adjust away part of the treatment effect under study. Similarly, clinical trial analyses may require adjustment for baseline confounders even if they do not improve prediction metrics. The m-XMT model allows for more intentional use of covariates to match scientific goals of analysis when compared to the RFRE.PO algorithm. The m-XMT (Clinical Input) model was the only appropriate model for assessing the effect of azithromycin presented in this manuscript.
The growing popularity of machine learning tools is likely to dominate analyses of high-dimensional data, and the RFRE.PO algorithm contributes to this literature as one of the first machine learning methods to be able to make dynamic predictions from recurrent event data. There are tuning parameters in random forest algorithms that may potentially be tuned for better performance, warranting further research. Some authors have studied parameter tuning (e.g., Scornet, 2018), but we have not seen work for longitudinal random forests to date. The development of (1) the theory of tuning parameters for longitudinal random forests and (2) the potential application of the censored longitudinal data structure for dynamic prediction to other machine learning algorithms are all important future research areas.
@@ Acknowledgment
The authors would like to thank the COPD Clinical Research Network, particularly our University of Michigan collaborator, Meilan Han, for the use of their data. Additional appreciation is given to the Editors, Associate Editor, and two anonymous reviewers whose comments significantly improved the presentation of the revised paper. The research is supported in part by a Precision Health Investigator Award from University of Michigan, Ann Arbor (AL, ZW). Conflict of Interest: None.